OpenStudy (anonymous):

What is cos(2a ) equal to if cos(a ) = 0.3?

OpenStudy (anonymous):

\[\cos(2a)=\cos^2(a)-\sin^2(a)\] use this formula and \[\sin^2 a+\cos^2 a=1\]

OpenStudy (mertsj):

If cos a = .3, a = 72.5 degrees and 2a =145 degrees. The cosine of 145 degrees is -.82

OpenStudy (anonymous):

how do we get 72.5

OpenStudy (mertsj):

Enter .3 into your calculator and push second function cosine.

OpenStudy (anonymous):

i dont see a function key on my calc

OpenStudy (mertsj):

Do you have a cosine button?

OpenStudy (anonymous):

yea

OpenStudy (mertsj):

What does it say right above the cosine button?

OpenStudy (anonymous):

you should click on \(cos^{-1}\) button

OpenStudy (anonymous):

cos^-1

OpenStudy (mertsj):

Bingo!!!

OpenStudy (anonymous):

and to do that you need go to 2nd floor

OpenStudy (anonymous):

kool do you know how to do this also

OpenStudy (anonymous):

there's button called 2nd

OpenStudy (anonymous):

yeah i know how to do that now...i was just wondering if you knew how to also do this question Give an algebraic expression for cos(sin^−1(x)). Here sin^−1(x) stands for the compositional inverse of sin(x).

OpenStudy (mertsj):

sin^-1(x) means "an angle whose sin is x" So if we call that angle y, for instance then we are talking about cos(y)

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

Or, if we are talking about the unit circle:

OpenStudy (mertsj):

Then x^2+y^2 = 1 and y = (1-x^2)^1/2 so the sin of the angle is x and the cosine is y or sqrt(1-x^2)